# On Bayesian Neural Networks with Dependent and Possibly Heavy-Tailed Weights

### v4

@@ 5 The infinite-width limit under the posterior
Theorem 5.1.
Assume conditions $(i)$ and $(ii)$ of Theorem 4.1 and
$\mathbb{P}_{{\rm prior}}(\mathrm{det}(\mathbb{K}^{(L+1)}(\mathbb{x}))>0)=1.$ (11)
@@
$\overrightarrow{\lim}(\mathbb{P}_{{\rm posterior}})_{Z^{(L+1)}(\mathbb{x})}=\mathcal{GM}_{n_{L+1}d}^{\mathbb{S}}(\lambda^{(L+1)}(\mathbb{x},\mathbb{y}),\boldsymbol{\Lambda}^{(L+1)}(\mathbb{x}))\quad\text{weakly}.$ (12)
Here $\mathbb{S}$ denotes the probability measure
@@
As clarified by Lemma 5.5, the assumption (11), which will be explored in the next section, is crucial only to identify the limiting law.
Remark 5.2.
Let $G^{(L+1)}(\mathbb{x})$ be the vector defined in (7) and suppose that it is distributed according to the Gaussian mixture law in (12). Then, under the probability measure $\mathbb{S}$, the rows of $\mathbb{G}^{(L+1)}(\mathbb{x})$ are dependent and, in general, not identical distributed. Indeed, under $\mathbb{S}$, the $i$-th row of $\mathbb{G}^{(L+1)}(\mathbb{x})$ is distributed according to a Gaussian mixture with random covariance $\mathbb{L}^{(L+1)}(\mathbb{x})^{-1}$ and random mean given by the $i$-th row of $2\mathbb{y}\mathbb{L}^{(L+1)}(\mathbb{x})^{-1}$; note that the rows of this latter matrix are all equal if and only if the rows of $\mathbb{y}$ are such.
Remark 5.3.
@@
Remark 5.4.
A simple modification of the proof of Theorem 5.1 shows that if $L\geq 2$, and we replace the condition $(i)$ of Theorem 4.1 with either the assumption (8) or the assumption (9), then, for any permutation $\pi$ of $[L]$, by Theorem 4.2 it follows that in (12) we can replace $\overrightarrow{\lim}$ with $\pi\overrightarrow{\lim}$.
The proof of Theorem 5.1 exploits the following Lemma 5.5 which, for the sake of completeness, will be proved at the end of this section.
Let $\{X_{n}\}_{n\geq 1}$ and $X$ be random variables defined of the probability space $(\Omega,\mathcal{F},\mathbb{P})$ and with values on some metric space. Consider on $(\Omega,\mathcal{F})$ the tilted probability measures:
$\mathrm{d}\mathbb{Q}_{n}\propto g(X_{n})\mathrm{d}\mathbb{P}\quad\text{and}\quad\mathrm{d}\mathbb{Q}\propto g(X)\mathrm{d}\mathbb{P}$
where $g$ is a bounded continuous function such that the normalizing constants are different from zero.
@@
Proof.
(Theorem 5.1) Thanks to Theorem 4.1 and Lemma 5.5, we only need to compute the law $\mathbb{Q}_{G^{(L+1)}(\mathbb{x})}$, where
$\mathrm{d}\mathbb{Q}\propto\exp\left(-\sum_{i=1}^{d}\|G^{(L+1)}(x(i))-y(i)\|^{2}\right)\mathrm{d}(\mathbb{P}_{{\rm prior}})_{G^{(L+1)}(\mathbb{x})}$
and
$(\mathbb{P}_{{\rm prior}})_{G^{(L+1)}(\mathbb{x})}=\mathcal{GM}_{n_{L+1}d}^{\mathbb{P}_{{\rm prior}}}(0,\mathbb{Id}_{n_{L+1}}\otimes\mathbb{K}^{(L+1)}(\mathbb{x})).$
Note that $\sum_{i=1}^{d}\|G^{(L+1)}(x(i))-y(i)\|^{2}=\|\mathbb{G}^{(L+1)}(\mathbb{x})-\mathbb{y}\|^{2}_{F}$. Let $\xi:=\mathrm{vec}(\boldsymbol{\xi}^{\top})$, where $\boldsymbol{\xi}:=[\xi(1)\ldots\xi(d)]$ and $\xi(i)\in\mathbb{R}^{n_{L+1}}$. By identity (2) ad the definition of Frobenius norm, we have
@@
$=\mathfrak{c}\,\mathbb{E}_{{\rm prior}}\Biggl[\left(\frac{1}{\det\mathbb{K}^{(L+1)}(\mathbb{x})}\right)^{\frac{n_{L+1}}{2}}\hskip-16.99998pt\exp\left(-\frac{1}{2}\mathrm{Tr}[2(\boldsymbol{\xi}-\mathbb{y})(\boldsymbol{\xi}-\mathbb{y})^{\top}-\boldsymbol{\xi}(\mathbb{K}^{(L+1)}(\mathbb{x}))^{-1}\boldsymbol{\xi}^{\top}]\right)\Biggr]\mathrm{d}\xi.$ (13)
Next, for ease of notation set $\lambda=\lambda^{(L+1)}(\mathbb{x},\mathbb{y})$, $\boldsymbol{\lambda}=\boldsymbol{\lambda}^{(L+1)}(\mathbb{x},\mathbb{y})$, $\boldsymbol{\Lambda}=\boldsymbol{\Lambda}^{(L+1)}(\mathbb{x})$, $\mathbb{L}=\mathbb{L}^{(L+1)}(\mathbb{x})$ and $\mathbb{K}=\mathbb{K}^{(L+1)}(\mathbb{x})$ and observe that by identity (2) we have
@@
$=\mathrm{Tr}((\boldsymbol{\xi}-\boldsymbol{\lambda})\mathbb{L}(\boldsymbol{\xi}-\boldsymbol{\lambda})^{\top}).$ (14)
On the other hand
$\mathrm{Tr}((\boldsymbol{\xi}-\boldsymbol{\lambda})\mathbb{L}(\boldsymbol{\xi}-\boldsymbol{\lambda})^{\top})=\mathrm{Tr}(\boldsymbol{\xi}\mathbb{L}\boldsymbol{\xi}^{\top})-2\mathrm{Tr}(\boldsymbol{\lambda}\mathbb{L}\boldsymbol{\xi}^{\top})+\mathrm{Tr}(\boldsymbol{\lambda}\mathbb{L}\boldsymbol{\lambda}^{\top}).$
By definition $\boldsymbol{\lambda}^{(L+1)}=2\mathbb{y}\mathbb{L}^{-1}$. Hence $\boldsymbol{\lambda}\mathbb{L}\boldsymbol{\xi}^{\top}=2\mathbb{y}\boldsymbol{\xi}^{\top}$. Plugging this expression in the previous identity and re-arranging in (14) yields
${\rm Tr}[\boldsymbol{\xi}\mathbb{L}\boldsymbol{\xi}^{\top}-4\mathbb{y}\boldsymbol{\xi}^{\top}]=(\xi-\lambda)^{\top}\boldsymbol{\Lambda}^{-1}(\xi-\lambda)-\mathrm{Tr}(\boldsymbol{\lambda}\mathbb{L}\boldsymbol{\lambda}^{\top}).$ (15)
@@
On combining this latter relation with (15) we have
${\rm Tr}[2(\boldsymbol{\xi}-\mathbb{y})(\boldsymbol{\xi}-\mathbb{y})^{\top}+\boldsymbol{\xi}\mathbb{K}^{-1}\boldsymbol{\xi}^{\top}]=(\xi-\lambda)^{\top}\boldsymbol{\Lambda}^{-1}(\xi-\lambda)-\mathrm{Tr}(\boldsymbol{\lambda}\mathbb{L}\boldsymbol{\lambda}^{\top})+2\mathrm{Tr}(\mathbb{y}\mathbb{y}^{\top}).$ (16)
Plugging the right-hand side of (16) in (14) yields
$\mathrm{d}\mathbb{Q}_{G^{(L+1)}(\mathbb{x})}(\xi)$
$=\mathfrak{c}\mathbb{E}_{{\rm prior}}\Biggl[(\mathrm{det}\mathbb{K}^{(L+1)}(\mathbb{x}))^{-n_{L+1}/2}\exp\left(\frac{1}{2}\mathrm{Tr}(\boldsymbol{\lambda}^{(L+1)}(\mathbb{x},\mathbb{y})\mathbb{L}^{(L+1)}(\mathbb{x})(\boldsymbol{\lambda}^{(L+1)}(\mathbb{x},\mathbb{y}))^{\top})\right)$
$\qquad\qquad\qquad\times\exp\left(-\frac{1}{2}(\xi-\lambda^{(L+1)}(\mathbb{x},\mathbb{y}))^{\top}(\boldsymbol{\Lambda}^{(L+1)}(\mathbb{x}))^{-1}(\xi-\lambda^{(L+1)}(\mathbb{x},\mathbb{y}))\right)\Biggr]\mathrm{d}\xi$
$=\mathfrak{c}\mathbb{E}_{{\rm prior}}\Biggl[\frac{(\mathrm{det}\mathbb{K}^{(L+1)}(\mathbb{x}))^{-n_{L+1}/2}\exp\left(\frac{1}{2}\mathrm{Tr}(\boldsymbol{\lambda}^{(L+1)}(\mathbb{x},\mathbb{y})\mathbb{L}^{(L+1)}(\mathbb{x})(\boldsymbol{\lambda}^{(L+1)}(\mathbb{x},\mathbb{y}))^{\top})\right)}{(2\pi)^{-(n_{L+1}d)/2}\mathrm{det}(\boldsymbol{\Lambda}^{(L+1)}(\mathbb{x}))^{-1/2}}$
$\qquad\times(2\pi)^{-(n_{L+1}d)/2}\mathrm{det}(\boldsymbol{\Lambda}^{(L+1)}(\mathbb{x}))^{-1/2}\exp\left(-\frac{1}{2}(\xi-\lambda^{(L+1)}(\mathbb{x},\mathbb{y}))^{\top}(\boldsymbol{\Lambda}^{(L+1)}(\mathbb{x}))^{-1}(\xi-\lambda^{(L+1)}(\mathbb{x},\mathbb{y}))\right)\Biggr]\mathrm{d}\xi$
$=\mathbb{E}_{\mathbb{S}}[\varphi_{(\lambda^{(L+1)}(\mathbb{x},\mathbb{y}),\boldsymbol{\Lambda}^{(L+1)}(\mathbb{x}))}(\xi)]\mathrm{d}\xi,$
@@
Proof.
(Lemma 5.5) Set
$\mathfrak{c}_{n}:=(\mathbb{E}[g(X_{n})])^{-1}\quad\text{and}\quad\mathfrak{c}:=(\mathbb{E}[g(X)])^{-1}.$
Let $S$ be the metric space where the random variables $X_{n}$, $n\geq 1$, and $X$ take values and let $f:S\to\mathbb{R}$ be a bounded and continuous function. We have
@@
Since $g(\cdot)$ and $f(\cdot)g(\cdot)$ are bounded and continuous functions by the weak convergence of $\mathbb{P}_{X_{n}}$ to $\mathbb{P}_{X}$ it follows
$\lim_{n\to\infty}\mathfrak{c}_{n}=\mathfrak{c}\quad\text{and}\quad\lim_{n\to\infty}I_{n}=0,$
and the proof is completed. ∎
6 Sufficient conditions for (11)
Condition (11) is crucial to identify the distribution of the infinite-width limit of the posterior output of the neural network (see Theorem 5.1). The following results provide sufficient conditions for (11).
Theorem 6.1.
Assume $\sigma$ Lipschitz continuous and non constant and
$a^{(\ell-1)}=0$ $\Rightarrow$ $\mathbb{P}_{{\rm prior}}(N_{\ell-1}((0,\infty))=\infty)=1$, $\forall$ $\ell\in\{2,\ldots,L+1\}$. (17)
@@ 6 Sufficient conditions for (11)
The data $x(1),\ldots,x(d)$ are linearly independent vectors, with $n_{0}\geq d$, (18)
then (11) holds.
Theorem 6.2.
Assume $\sigma$ Lipschitz continuous and nonlinear and
$a^{(1)}>0$ and $a^{(\ell-1)}=0$ $\Rightarrow$ $\mathbb{P}_{{\rm prior}}(N_{\ell-1}((0,\infty))=\infty)=1$, $\forall$ $\ell\in\{3,\ldots,L+1\}$. (19)
@@
If $C_{B}>0$ then the data $x(1),\ldots,x(d)$ are all distinct (20)
@@
If $C_{B}=0$ then the data $x(1),\ldots,x(d)$ are pairwise non-proportional, (21)
then (11) holds.
@@
Proposition 6.3.
Assume $\sigma$ Lipschitz continuous and non constant, and let $\ell\in\{2,\ldots,L+1\}$ be fixed. If
$\mathbb{P}_{{\rm prior}}(\mathrm{det}(\mathbb{K}^{(\ell-1)}(\mathbb{x}))>0)=1$ (22)
@@
$a^{(\ell-1)}=0$ $\Rightarrow$ $\mathbb{P}_{{\rm prior}}(N_{\ell-1}((0,\infty))=\infty)=1$, (23)
@@
$\mathbb{P}_{{\rm prior}}(\mathrm{det}(\mathbb{K}^{(\ell)}(\mathbb{x}))>0)=1.$ (24)
The proof of Proposition 6.3 exploits the following results, which are proved at the end of the section.
Proposition 6.4.
Let $\sigma$ be Lipschitz continuous and non constant, and let $X$ be a $d$-dimensional random vector with a density with respect to the Lebesgue measure which is strictly positive on $\mathbb{R}^{d}$. Then there exist $A=A_{\sigma}\in\mathcal{B}(\mathbb{R}^{d})$ (AA does not depend on $X$) and $\varphi_{\sigma(X)}:A\to[0,\infty)$ measurable such that
$\mathbb{P}(\sigma(X)\in C)=\int_{C}\varphi_{\sigma(X)}(y)\mathrm{d}y$, for any $C\in\mathcal{B}(A)$, and $\mathbb{P}(\sigma(X)\in A)>0$. (25)
Proposition 6.5.
Let $X$ be a real-valued $m$-dimensional (column) random vector and suppose that its law and the Lebesgue measure on $\mathbb{R}^{m}$ are not singular and that $\mathbb{E}[XX^{\top}]<\infty$ (i.e., all the entries of the matrix $\mathbb{E}[XX^{\top}]$ are finite). Then the matrix $\mathbb{E}[XX^{\top}]$ is positive definite.
Proposition 6.6.
Let $X_{1},\ldots,X_{r}$ be $r$ (column) random vectors with values in $\mathbb{R}^{d}$, $1\leq r\leq d$. If the law of ${\rm vec}((X_{1},\ldots,X_{r}))$ is absolutely continuous with respect to the Lebesgue measure on $\mathbb{R}^{dr}$, then the random vectors $\{X_{1},\ldots,X_{r}\}$ are linearly independent almost surely.
6.1 Proof of Theorems 6.1 and 6.2
Proof of Theorem 6.1.
The assumption (18) implies $\mathrm{det}(\mathbb{K}^{(1)}(\mathbb{x}))>0$ (indeed, it is easily verified that if the columns of a matrix $\mathbb{A}$ are linearly independent then the Gram matrix $\mathbb{A}^{\top}\mathbb{A}$ is positive definite). The claim follows by Proposition 6.3. ∎
Proof of Theorem 6.2.
By Theorems 6 and 7 in [8] we have that the matrix $\mathbb{E}_{{\rm prior}}[\sigma(\zeta_{1}^{(1)}(\mathbb{x}))\sigma(\zeta_{1}^{(1)}(\mathbb{x}))^{\top}]$ is positive definite. Recall that
$\mathbb{K}^{(2)}(\mathbb{x})=C_{B}1_{d}1_{d}^{\top}+C_{W}\Biggl(a^{(1)}\mathbb{E}_{{\rm prior}}[\sigma(\zeta_{1}^{(1)}(\mathbb{x}))\sigma(\zeta_{1}^{(1)}(\mathbb{x}))^{\top}]+\sum_{j=1}^{N_{1}((0,\infty))}T_{j}^{(1)}\sigma(\zeta_{j}^{(1)}(\mathbb{x}))\sigma(\zeta_{j}^{(1)}(\mathbb{x}))^{\top}\Biggr).$

### v5

@@ 5. The infinite-width limit under the posterior
Theorem 3.
Assume conditions $(i)$ and $(ii)$ of Theorem 1 and
(5.2) $\mathbb{P}_{{\rm prior}}(\mathrm{det}(\mathbb{K}^{(L+1)}(\mathbb{x}))>0)=1.$
@@
(5.3) $\lim_{n_{L}\to\infty}\ldots\lim_{n_{1}\to\infty}(\mathbb{P}_{{\rm posterior}})_{Z^{(L+1)}(\mathbb{x})}=\mathcal{GM}_{n_{L+1}d}^{\mathbb{S}}(\lambda^{(L+1)}(\mathbb{x},\mathbb{y}),\boldsymbol{\Lambda}^{(L+1)}(\mathbb{x}))\quad\text{weakly.}$
Here $\mathbb{S}$ denotes the probability measure on $(\Omega,\mathcal{F})$ defined by
@@
As clarified by Lemma 2, the assumption (5.2), which will be explored in the next section, is crucial only to identify the limiting law.
Remark 1.
Let $G^{(L+1)}(\mathbb{x})$ be the r.v. defined in (4.4) and suppose that it is distributed according to the Gaussian mixture law in (5.3). Then, under the probability measure $\mathbb{S}$, the rows of $\mathbb{G}^{(L+1)}(\mathbb{x})$ are dependent and, in general, not identical distributed. Indeed, under $\mathbb{S}$, the $i$-th row of $\mathbb{G}^{(L+1)}(\mathbb{x})$ is distributed according to a Gaussian mixture with random covariance $\mathbb{L}^{(L+1)}(\mathbb{x})^{-1}$ and random mean given by the $i$-th row of $2\mathbb{y}\mathbb{L}^{(L+1)}(\mathbb{x})^{-1}$; note that the rows of this latter matrix are all equal if and only if the rows of $\mathbb{y}$ are such.
Remark 2.
@@
Remark 3.
A simple modification of the proof of Theorem 3 shows that if $L\geq 2$, and we replace the condition $(i)$ of Theorem 1 with either the assumption (4.5) or the assumption (4.6), then by Theorem 2 it follows that in (5.3) we can replace $\lim_{n_{L}\to\infty}\ldots\lim_{n_{1}\to\infty}$ with $\pi\lim$.
The proof of Theorem 3 exploits the following Lemma 2 which, for the sake of completeness, will be proved at the end of this section.
Let $\{X_{n}\}_{n\geq 1}$ and $X$ be r.v.’s defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$ and with values on some metric space. Consider on $(\Omega,\mathcal{F})$ the tilted probability measures: $\mathrm{d}\mathbb{Q}_{n}\propto g(X_{n})\mathrm{d}\mathbb{P}$ and $\mathrm{d}\mathbb{Q}\propto g(X)\mathrm{d}\mathbb{P}$ where $g$ is a bounded continuous function such that the normalizing constants are different from zero.
@@
Proof of Theorem 3.
Thanks to Theorem 1 and Lemma 2, we only need to compute the law $\mathbb{Q}_{G^{(L+1)}(\mathbb{x})}$, where
$\mathrm{d}\mathbb{Q}\propto\exp\left(-\sum_{i=1}^{d}\|G^{(L+1)}(x(i))-y(i)\|^{2}\right)\mathrm{d}\mathbb{P}_{{\rm prior}}$
and $(\mathbb{P}_{{\rm prior}})_{G^{(L+1)}(\mathbb{x})}=\mathcal{GM}_{n_{L+1}d}^{\mathbb{P}_{{\rm prior}}}(0,\mathbb{Id}_{n_{L+1}}\otimes\mathbb{K}^{(L+1)}(\mathbb{x})).$ Note that $\sum_{i=1}^{d}\|G^{(L+1)}(x(i))-y(i)\|^{2}=\|\mathbb{G}^{(L+1)}(\mathbb{x})-\mathbb{y}\|^{2}_{F}$. Let $\xi:=\mathrm{vec}(\boldsymbol{\xi}^{\top})$, where $\boldsymbol{\xi}:=[\xi(1)\ldots\xi(d)]$ and $\xi(i)\in\mathbb{R}^{n_{L+1}}$. By identity (3.2) ad the definition of Frobenius norm, we have
@@
(5.4) $=\mathfrak{c}\,\mathbb{E}_{{\rm prior}}\Biggl[\left(\frac{1}{\det\mathbb{K}^{(L+1)}(\mathbb{x})}\right)^{\frac{n_{L+1}}{2}}\hskip-16.99998pt\exp\left(-\frac{1}{2}\mathrm{Tr}[2(\boldsymbol{\xi}-\mathbb{y})(\boldsymbol{\xi}-\mathbb{y})^{\top}-\boldsymbol{\xi}(\mathbb{K}^{(L+1)}(\mathbb{x}))^{-1}\boldsymbol{\xi}^{\top}]\right)\Biggr]\mathrm{d}\xi.$
Next, for ease of notation set $\lambda=\lambda^{(L+1)}(\mathbb{x},\mathbb{y})$, $\boldsymbol{\lambda}=\boldsymbol{\lambda}^{(L+1)}(\mathbb{x},\mathbb{y})$, $\boldsymbol{\Lambda}=\boldsymbol{\Lambda}^{(L+1)}(\mathbb{x})$, $\mathbb{L}=\mathbb{L}^{(L+1)}(\mathbb{x})$ and $\mathbb{K}=\mathbb{K}^{(L+1)}(\mathbb{x})$ and observe that by identity (3.2) we have
@@
(5.5) $=\mathrm{Tr}((\boldsymbol{\xi}-\boldsymbol{\lambda})\mathbb{L}(\boldsymbol{\xi}-\boldsymbol{\lambda})^{\top}).$
On the other hand $\mathrm{Tr}((\boldsymbol{\xi}-\boldsymbol{\lambda})\mathbb{L}(\boldsymbol{\xi}-\boldsymbol{\lambda})^{\top})=\mathrm{Tr}(\boldsymbol{\xi}\mathbb{L}\boldsymbol{\xi}^{\top})-2\mathrm{Tr}(\boldsymbol{\lambda}\mathbb{L}\boldsymbol{\xi}^{\top})+\mathrm{Tr}(\boldsymbol{\lambda}\mathbb{L}\boldsymbol{\lambda}^{\top}).$ By definition $\boldsymbol{\lambda}^{(L+1)}=2\mathbb{y}\mathbb{L}^{-1}$. Hence $\boldsymbol{\lambda}\mathbb{L}\boldsymbol{\xi}^{\top}=2\mathbb{y}\boldsymbol{\xi}^{\top}$. Plugging this expression in the previous identity and re-arranging in (5.5) yields
(5.6) ${\rm Tr}[\boldsymbol{\xi}\mathbb{L}\boldsymbol{\xi}^{\top}-4\mathbb{y}\boldsymbol{\xi}^{\top}]=(\xi-\lambda)^{\top}\boldsymbol{\Lambda}^{-1}(\xi-\lambda)-\mathrm{Tr}(\boldsymbol{\lambda}\mathbb{L}\boldsymbol{\lambda}^{\top}).$
@@
On combining this latter relation with (5.6) we have
(5.7) ${\rm Tr}[2(\boldsymbol{\xi}-\mathbb{y})(\boldsymbol{\xi}-\mathbb{y})^{\top}+\boldsymbol{\xi}\mathbb{K}^{-1}\boldsymbol{\xi}^{\top}]=(\xi-\lambda)^{\top}\boldsymbol{\Lambda}^{-1}(\xi-\lambda)-\mathrm{Tr}(\boldsymbol{\lambda}\mathbb{L}\boldsymbol{\lambda}^{\top})+2\mathrm{Tr}(\mathbb{y}\mathbb{y}^{\top}).$
Plugging the right-hand side of (5.7) in (5.5) yields
$\mathrm{d}\mathbb{Q}_{G^{(L+1)}(\mathbb{x})}(\xi)$ $=\mathfrak{c}\mathbb{E}_{{\rm prior}}\Biggl[(\mathrm{det}\mathbb{K}^{(L+1)})^{-n_{L+1}/2}\exp\left(\frac{1}{2}\mathrm{Tr}(\boldsymbol{\lambda}^{(L+1)}\mathbb{L}^{(L+1)}(\boldsymbol{\lambda}^{(L+1)})^{\top})\right)$
$\qquad\qquad\qquad\times\exp\left(-\frac{1}{2}(\xi-\lambda^{(L+1)})^{\top}(\boldsymbol{\Lambda}^{(L+1)})^{-1}(\xi-\lambda^{(L+1)})\right)\Biggr]\mathrm{d}\xi$
$=\mathfrak{c}\mathbb{E}_{{\rm prior}}\Biggl[\frac{(\mathrm{det}\mathbb{K}^{(L+1)})^{-n_{L+1}/2}\exp\left(\frac{1}{2}\mathrm{Tr}(\boldsymbol{\lambda}^{(L+1)}\mathbb{L}^{(L+1)}(\boldsymbol{\lambda}^{(L+1)})^{\top})\right)}{(2\pi)^{-(n_{L+1}d)/2}\mathrm{det}(\boldsymbol{\Lambda}^{(L+1)})^{-1/2}}$
$\qquad\qquad\times\mathrm{det}(\boldsymbol{\Lambda}^{(L+1)})^{-1/2}\exp\left(-\frac{1}{2}(\xi-\lambda^{(L+1)})^{\top}(\boldsymbol{\Lambda}^{(L+1)})^{-1}(\xi-\lambda^{(L+1)})\right)\Biggr]\mathrm{d}\xi$
$=\mathbb{E}_{\mathbb{S}}[\varphi_{(\lambda^{(L+1)},\boldsymbol{\Lambda}^{(L+1)})}(\xi)]\mathrm{d}\xi,$
@@
Proof of Lemma 2.
Set $\mathfrak{c}_{n}:=(\mathbb{E}[g(X_{n})])^{-1}$ and $\mathfrak{c}:=(\mathbb{E}[g(X)])^{-1}.$ Let $S$ be the metric space where the r.v.’s $X_{n}$, $n\geq 1$, and $X$ take values and let $f:S\to\mathbb{R}$ be a bounded and continuous function. We have
@@
Since $g(\cdot)$ and $f(\cdot)g(\cdot)$ are bounded and continuous functions by the weak convergence of $\mathbb{P}_{X_{n}}$ to $\mathbb{P}_{X}$ it follows $\lim_{n\to\infty}\mathfrak{c}_{n}=\mathfrak{c}$ and $\lim_{n\to\infty}I_{n}=0$, and the proof is completed. ∎
6. Sufficient conditions for (5.2)
Condition (5.2) is crucial to identify the distribution of the infinite-width limit of the posterior output of the neural network (see Theorem 3). The following results provide sufficient conditions for (5.2).
Theorem 4.
Let the notation and assumption $(ii)$ of Theorem 1 prevail. Assume $\sigma$ Lipschitz continuous and non constant and
(6.1) $a^{(\ell-1)}=0$ $\Rightarrow$ $\mathbb{P}_{{\rm prior}}(N_{\ell-1}((0,\infty))=\infty)=1$, $\forall$ $\ell\in\{2,\ldots,L+1\}$.
@@
(6.2) The data $x(1),\ldots,x(d)$ are linearly independent vectors, with $n_{0}\geq d$,
then (5.2) holds.
Theorem 5.
Let the notation and assumption $(ii)$ of Theorem 1 prevail. Assume $\sigma$ Lipschitz continuous and nonlinear and
(6.3) $a^{(1)}>0$ and $a^{(\ell-1)}=0$ $\Rightarrow$ $\mathbb{P}_{{\rm prior}}(N_{\ell-1}((0,\infty))=\infty)=1$, $\forall$ $\ell\in\{3,\ldots,L+1\}$.
@@
(6.4) If $C_{B}>0$ then the data $x(1),\ldots,x(d)$ are all distinct
@@
(6.5) If $C_{B}=0$ then the data $x(1),\ldots,x(d)$ are pairwise non-proportional,
then (5.2) holds.
@@
Proposition 1.
Let the notation and assumption $(ii)$ of Theorem 1 prevail, and assume $\sigma$ Lipschitz continuous and non constant. If for $\ell\in\{2,\ldots,L+1\}$ it holds
(6.6) $\mathbb{P}_{{\rm prior}}(\mathrm{det}(\mathbb{K}^{(\ell-1)}(\mathbb{x}))>0)=1$
@@
(6.7) $a^{(\ell-1)}=0$ $\Rightarrow$ $\mathbb{P}_{{\rm prior}}(N_{\ell-1}((0,\infty))=\infty)=1$,
@@
(6.8) $\mathbb{P}_{{\rm prior}}(\mathrm{det}(\mathbb{K}^{(\ell)}(\mathbb{x}))>0)=1.$
The proof of Proposition 1 exploits the following results, which are proved at the end of the section.
Proposition 2.
Let $\sigma$ be Lipschitz continuous and non constant, and let $X$ be a $d$-dimensional r.v. with a density with respect to the Lebesgue measure which is strictly positive on $\mathbb{R}^{d}$. Then there exist $A=A_{\sigma}\in\mathcal{B}(\mathbb{R}^{d})$ (AA does not depend on $X$) and $\varphi_{\sigma(X)}:A\to[0,\infty)$ measurable such that
(6.9) $\mathbb{P}(\sigma(X)\in C)=\int_{C}\varphi_{\sigma(X)}(y)\mathrm{d}y$, for any $C\in\mathcal{B}(A)$, and $\mathbb{P}(\sigma(X)\in A)>0$.
Proposition 3.
Let $X$ be a real-valued $m$-dimensional (column) r.v. and suppose that its law and the Lebesgue measure on $\mathbb{R}^{m}$ are not singular and that $\mathbb{E}[XX^{\top}]<\infty$ (i.e., all the entries of the matrix $\mathbb{E}[XX^{\top}]$ are finite). Then the matrix $\mathbb{E}[XX^{\top}]$ is positive definite.
Proposition 4.
Let $X_{1},\ldots,X_{r}$ be $r$ (column) r.v.’s with values in $\mathbb{R}^{q}$, $1\leq r\leq q$. If the law of ${\rm vec}((X_{1},\ldots,X_{r}))$ is absolutely continuous with respect to the Lebesgue measure on $\mathbb{R}^{qr}$, then the random vectors $\{X_{1},\ldots,X_{r}\}$ are linearly independent almost surely.
6.1. Proof of Theorems 4 and 5
Proof of Theorem 4.
The assumption (6.2) implies $\mathrm{det}(\mathbb{K}^{(1)}(\mathbb{x}))>0$ (indeed, it is easily verified that if the columns of a matrix $\mathbb{A}$ are linearly independent then the Gram matrix $\mathbb{A}^{\top}\mathbb{A}$ is positive definite). The claim follows by Proposition 1. ∎
Proof of Theorem 5.
By Theorems 6 and 7 in [8] we have that the matrix
$\mathbb{E}_{{\rm prior}}[\sigma(\zeta_{1}^{(1)}(\mathbb{x}))\sigma(\zeta_{1}^{(1)}(\mathbb{x}))^{\top}]$
is positive definite. Recall that
$\mathbb{K}^{(2)}(\mathbb{x})$
$=C_{B}1_{d}1_{d}^{\top}+C_{W}\Biggl(a^{(1)}\mathbb{E}_{{\rm prior}}[\sigma(\zeta_{1}^{(1)}(\mathbb{x}))\sigma(\zeta_{1}^{(1)}(\mathbb{x}))^{\top}]+\sum_{j=1}^{N_{1}((0,\infty))}T_{j}^{(1)}\sigma(\zeta_{j}^{(1)}(\mathbb{x}))\sigma(\zeta_{j}^{(1)}(\mathbb{x}))^{\top}\Biggr).$